OpenStudy (loser66):

Show that \(\dfrac{3^n}{n!}\) converges to 0 Please help.

OpenStudy (loser66):

@dan815

OpenStudy (zarkon):

\[\frac{3}{n}\frac{3}{n-1}\cdots\frac{3}{3}\frac{3}{2}\frac{3}{1}=\frac{3}{n}\times\frac{3}{n-1}\cdots\frac{3}{3}\times\frac{3}{2}\frac{3}{1}\le\frac{27}{2n}\]

OpenStudy (zarkon):

note \[\frac{3}{n-1}\cdots\frac{3}{3}\le 1\]

OpenStudy (loser66):

Oh, You try to put in sandwich theorem, right?

OpenStudy (zarkon):

yes...\[0\le\frac{3^n}{n!}\le\frac{27}{2n}\]

OpenStudy (loser66):

the far right is = (27/2) (1/n) which goes to 0 when n goes to infinitive, right?

OpenStudy (zarkon):

yes

OpenStudy (loser66):

Is there any logic between the line? That is when we assume that the middle term has limit, it will be 0 .

OpenStudy (zarkon):

by the squeeze theorem it has to be zero

OpenStudy (zarkon):

it is trapped between to things that are going to zero

OpenStudy (loser66):

I lost my credits because of this mistake many times, but still not know how to fix it. My Prof commented : "How can you know whether the limit of the term exist or not?"

OpenStudy (zarkon):

it is given in the theorem... http://en.wikipedia.org/wiki/Squeeze_theorem#Statement

OpenStudy (loser66):

He said : "If it exists, then it is 0, but we don't know whether it exists or not yet"

OpenStudy (zarkon):

squeeze theorem tells us it exists

OpenStudy (zarkon):

and it has to have the same value as the other limits

OpenStudy (loser66):

Thank you. I review for final and I do not want to make mistakes. :)

OpenStudy (loser66):

@Zarkon This is the way how I lost my credit :(

OpenStudy (zarkon):

ic what you did

OpenStudy (zarkon):

state the inequalities without limits like I did above... then find the limits of the left and right function.... finally (if the limits are the same) state by the squeeze theorem that the limit exists and is equal to the previous limits

OpenStudy (zarkon):

understand?

OpenStudy (loser66):

Yes, Sir. Thank you. :)

OpenStudy (dan815):

harsh marker a complete 0 :P